OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation x - 4y = 4.

OpenStudy (anonymous):

The equation is currently in standard form. You have to get it into slope-intercept form. \[x - 4y = 4\]\[-4y = -x + 4\]\[y = \frac{-x + 4}{-4}\]\[y = \frac{x}{4} - 1\]\[y = \frac{1}{4}x - 1\]The slope-intercept form is y = mx + b where m = slope and b = y-intercept. What are the slope and y-intercept using this knowledge?

OpenStudy (anonymous):

First change that equation into the standard y=mx +b form (where m is the slope, x and y are the coordinates (x,y) and b is the y intercept) You will have to substract x from both sides and get -4y = 4-x Then divide both sides by -4 get y by itself, getting you y= (-x+4)/-4 y=(1\4)x -1 The slope is (1/4, which is m) The y-intercept is (-1, which is b) Complete. Simple enough. :)

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

np :)